In a small US town of 1,000 people there are 100 Children under 16, 100 retirees, 600, people with full-time jobs, 30 people with part-time jobs, 30 non-working full-time students aged 16 and over, and 40 full-time homemakers. the remaining people do not have jobs but want them. All except ten of these had actively searched for a job in the previous four weeks of the time when the Census Bureau interviewed in this town.

  1. Based on the data above, calculate the size of the town’s labor force.
  2. Based on the data above, calculate the town’s unemployment rate.
Answer #1

(a)

Labor force is consisted of both employed and unemployed.

Employed includes both full time employed and part time employed.

Employed = Full time employed + Part time employed

Employed = 600 + 30 = 630

The number of employed is 630 people.

People without job = Total population - Children under 16 - Retirees - Full time employed - Part time employed - Full time students - Full time homemakers

People without jobs = 1,000 - 100 - 100 - 600 - 30 - 30 - 40

People without jobs = 1,000 - 900 = 100

The people without jobs is 100 persons.

Out of these 100 persons, 10 workers has not actively searched for job in previous four weeks.

So, these 10 are discouraged workers and are not counted as employed.

Thus,

the number of unemployed is 90 persons.

Calculate the labor force -

Labor force = Employed + Unemployed = 630 + 90 = 720

The size of the town's labor force is 720 people.

(b)

Calculate the unemployment rate -

Unemployment rate = [Unemployed/Labor force] * 100

Unemployment rate = [90/720] * 100 = 12.5%

Thus,

The unemployment rate is 12.5%